OpenStudy (anonymous):

PLEASE HELP!!!! A company has found that the relationship between the price p and the demand x for a particular product is given approximately by p=1269−0.2x2. The company also knows that the cost of producing the product is given by C(x)=830+398x. (A) Find P(x), the profit function. P(x) = Now use the profit function to do the following: (B) Find the average of the x values of all local maxima of P. Note: If there are no local maxima, enter "none". Average of x values =

OpenStudy (perl):

profit = revenue - cost

OpenStudy (perl):

Profit = revenue - cost we write this as function notation: P(x) = R(x) - C(x) P(x) = (number of units demanded x )* (price per unit when demand is x)- cost of making x units P(X) = x*p - C(x) P(X) = x * (1269 - .2x^2) - (830 + 398x)

OpenStudy (anonymous):

Thank you, how exactly do you find part b?

OpenStudy (perl):

first find the local maxima of P, then average them

OpenStudy (perl):

this is P(x)

OpenStudy (perl):

these are the x values of the local maxima http://prntscr.com/6xh8gx

OpenStudy (perl):

the average is right in between , and due to symmetry, the average must be zero

OpenStudy (anonymous):

It's saying zero is incorrect

OpenStudy (perl):

is the first part correct

OpenStudy (anonymous):

it wanted the exact amount at 38.1007, thank you very much

OpenStudy (perl):

the average is 38.10007?

OpenStudy (anonymous):

i believe it just wanted the local maxima

OpenStudy (perl):

these are x values of the local maximum. and it is 38.1007, -38.1007 http://www.wolframalpha.com/input/?i=critical+points+x+*+%281269+-+.2x^2%29+-+%28830+%2B+398x%29

OpenStudy (perl):

so that sounds right :)